OpenStudy (anonymous):

Find the following quotient: x^2-4 ------- x+2

OpenStudy (anonymous):

= (x+2)(x-2) / (x+2) = x-2

OpenStudy (anonymous):

\[x^2-y^2=(x+y)(x-y)\] \[\frac{(x)^2-(2)^2}{x+2}\] \[=\frac{\cancel{(x+2)}(x-2)}{\cancel{(x+2)}}\] \[=x-2\]